What mass of phosphorus pentachloride would be produced by the complete reaction of 6.66 g of P4? Report your answer to the hundredths of a gram without units.
Answer #1

Molar mass of P4 = 123.88 g/mol

mass of P4 = 6.66 g

mol of P4 = (mass)/(molar mass)

= 6.66/1.239*10^2

= 5.376*10^-2 mol

Balanced chemical equation is:

P4 + 10 Cl2 -> 4 PCl5

According to balanced equation

mol of PCl5 formed = (4/1)* moles of P4

= (4/1)*5.376*10^-2

= 0.215 mol

Molar mass of PCl5,

MM = 1*MM(P) + 5*MM(Cl)

= 1*30.97 + 5*35.45

= 208.22 g/mol

mass of PCl5 = number of mol * molar mass

= 0.215*2.082*10^2

= 44.78 g

Answer: 44.78 g